OpenStudy (anonymous):

The function f(x) = x3 + 4 is transformed to give a new function, g(x). The graph of g(x) [red] shows a horizontal shrink compared to the graph of f(x) [blue]. What is the function g(x)?

OpenStudy (anonymous):

A.)g(x)= 2(x^3 + 4) B.) g(x)= 1/2 (x^3 + 4) C.) g(x)= 2x^3 + 4 D.) g(x)= 1/2 x^3 + 4

OpenStudy (anonymous):

\[A. g(x)=2(x^3+4) \] B.) \[g(x)=\frac{ 1 }{ 2 }(x^3+4)\] C). \[g(x)=2x^3+4\] D.) \[g(x)=\frac{ 1 }{ 2 } x^3+4\]